Ch 24 Homework x + -ox W. € → C newconnect.mheducation.com/flow/connect.html ☆ Suggested Sites Imported From IE Ch 24 HomeworCompute the payback period for each of these two separate investments: A new operating system for an existing machine is expected to cost $270,000 and have a useful life of five years. The system yields an incremental after-tax income of $77,884 each year after deducting its straight-line depreciation. The predicted salvage value of the system is $10,000. A machine costs $180,000, has a $14,000 salvage value, is expected to last nine years, and will generate an after-tax income of $43,000 per year after straight-line depreciation.
Answer #1

Solution:

a. Annual depreciation = (270000- 10000) / 5 = $52000

Annual cash flows = 77884 + 52000 = $129,884

b. Annual depreciation = (180000- 14000) / 9 = $18,444

Annual cash flows = 43000 + 18444 = $61,444

Payback Period
Choose Numerator / Choose Denominator = Payback Period
Cost of investment / Annual net Cash flow = Payback Period
$2,70,000 / $1,29,884 = 2.08
$1,80,000 / $61,444 = 2.93